# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v27

@@ Abstract
Keywords: Contingent Claims, Price Discovery, Options, Straddle, Volatility Smile.
@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information about an asset’s expected payoff, that information is incorporated into the asset price. More broadly, agents often have information about an asset’s payoff beyond its mean, such as its volatility, which requires trading derivatives. Indeed, a principal utility of derivatives is to offer investors targeted exposure to various aspects of the underlying asset’s payoff distribution. Such derivative trading has long been standard practice, as covered in popular textbooks such as Hull (2015). This means that various information beyond the expected payoff is incorporated into derivatives prices.
Yet, despite their practical significance, the existing literature lacks a comprehensive model that articulates how derivative markets systematically incorporate diverse information signals into prices. This gap arises largely due to the challenges of modeling the complex information dynamics across derivatives. We address this gap by developing a model that fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. Integrating the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985), we obtain general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
@@
Replicating derivatives, such as options, using AD securities yields equivalent formulations of our model that account for the structural linkages across derivatives. In an equivalent options formulation, the informed trader has information about an underlying asset’s payoff distribution and trades options, with no restriction on the payoff distribution or his information.
@@
The informed demand formula we obtain is simple and practical. It constructs the informed trader’s contingent-claim portfolio by assigning positive portfolio weight to the true payoff distribution and negative weights to the others. Translated to options, the formula prescribes option strategies for trading on any aspect of the underlying payoff distribution. In particular, it rationalizes empirically observed option trading strategies as equilibrium outcomes—the first paper to do so.
@@
When only the underlying asset is traded, our model reduces to the single-asset price discovery framework of Kyle (1985), where the equilibrium supply function is linear. The slope of the inverse supply function—known as “Kyle’s lambda”—determines the price impact and the extent to which demand reveals information about the asset’s payoff. This slope is proportional to the variation of the asset’s payoff. In the more general AD setting, demand for one security impacts prices across all markets based on the information it conveys about other securities. Building on the intuition behind Kyle’s lambda, we show that the equilibrium price impact between two securities is proportional to the covariance of their payoffs. This generalizes Kyle’s lambda to an infinite-dimensional matrix of payoff covariances.
@@
In option markets, trades in one option typically affect the prices of others. This occurs because trades in options jointly convey information about the underlying payoff distribution. For example, a common volatility trade comprises a put-call options pair. Thus, trades in such a pair jointly convey information about volatility. Our general characterization of such cross-market spillovers introduces new empirical measures and predictions that link option cross-strike dynamics to different aspects of the underlying payoff. Moreover, we offer an intuitive equilibrium explanation for the “volatility smile,” an empirical regularity in option-implied volatilities that has been widely studied but limited to reduced-form models.
@@
The joint supply schedules for all AD securities form a pricing kernel for state-contingent claims under asymmetric information. In contrast to the one-dimensional linear supply function in the single-asset case, the pricing kernel is a high-dimensional nonlinear object that prices all derivatives simultaneously. While the pricing of derivatives has been extensively studied from the hedging perspective since Black and Scholes (1973), we extend beyond the hedging perspective to asymmetric information, providing a microfoundation for those trading and price patterns unexplained by hedging motives alone.
@@
Empirical evidence (e.g., Roll et al. (2010)) suggests derivatives improve market efficiency, but theoretical explanations are limited. Our model offers a clear explanation: the linear equilibrium of Kyle (1985) is not robust with respect to the introduction of derivative securities, as the absence of derivative markets constrains the informed trader to suboptimal linear demands. Derivatives allow the informed trader to optimally reduce his demand for low-probability states (see Example 7.1). By incorporating information that cannot be reflected in linear trades of the underlying asset, derivatives improve market efficiency.
@@
In a complete market with derivatives, we show that the information efficiency of all prices is summarized in a single measure. Notably, this measure does not depend on the possible payoff distributions or noise trading levels across markets. Thus, the informativeness of derivatives prices is robust to specific information details.
@@
Kyle (1985) is the seminal model for single-asset price discovery. Extensions by Caballe and Krishnan (1994), Foster and Viswanathan (1996), and Back et al. (2000) consider settings with multiple insiders holding differential information across assets. Rochet and Vila (1994) relax the assumption of normal payoff distributions. We expand the scope of this literature to encompass joint price discovery across contingent claim markets.
@@
Empirical research has consistently shown that trades and prices across options markets reflect information beyond the expected payoff. Pan and Poteshman (2006) and Ni et al. (2008) find that option trading volumes predict future stock returns and volatility, while Goyal and Saretto (2009) show that volatility mispricing can be exploited through option straddles. Other studies show that option order flows can signal firm value and market expectations of future volatility—see Roll et al. (2009), Augustin et al. (2019), and Cao et al. (2005). Despite this empirical evidence for joint price formation across option markets, theoretical models of informed option trading have been limited to no-option vs. single-option comparisons, as in Back (1993), Biais and Hillion (1994), Easley et al. (1998), and Collin-Dufresne et al. (2021). This creates a theoretical gap in understanding how trades and prices across options jointly reflect broader information. Our model helps bridge this gap; it provides a unified theoretical framework that consolidates empirical findings and suggests new avenues of empirical research.
@@
There have also been theoretical studies on non-strategic informed trading across multiple securities, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). Non-strategic models do not explain the common derivatives trading strategies or observed cross-strike empirical regularities; moreover, Admati (1985) and Chabakauri et al. (2022) impose parametric assumptions on the payoff distribution while we do not. Additionally, we derive empirical predictions regarding how different payoff information (e.g., higher moments ) can be inferred from price impact across options. This new insight is unique to the strategic setting.
@@
Many option pricing and financial econometric models have sought to capture the volatility smile by correcting misspecifications in the Black and Scholes (1973) framework. Examples include the jump-diffusion models of Merton (1976)-type, [^1] See Bates (1991). the stochastic volatility models of Heston (1993)-type, [^2] See Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), and Aït-Sahalia et al. (2021). and the local volatility models of Dupire (1994)-type. [^3] See Berestycki et al. (2002) and Carr and Cousot (2012). Our framework moves beyond reduced-form model misspecifications and directly explains the strike structure of implied volatility via cross-strike information dynamics.
@@
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results in a simplified setting. Section 3 presents the general model. Section 4 shows that the uninformed trader’s cross-market inferences are summarized by an intuitive sufficient statistic. Section 5 analyzes the informed trader’s cross-market price impact by decomposing his first-order condition. Section 6 establishes equilibrium through a canonical reduction of the trading game. While this canonical reduction is a key modeling innovation that makes our general model tractable, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be understood from Example 6.7 without delving into the precise details. Section 7 discusses the characterization of price discovery for contingent claims—the informed demand, price impact, and information efficiency of prices. Section 8 explores the immediate empirical implications of our results for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@ 2 Basic Intuition
We begin with a simplified setting that encapsulates the basic economic intuition we bring out. In the general model, we will develop the robustness of this intuition—without restriction on the number of securities, payoff distributions, or types of private information.
@@ 3 General Model
As in Section 2, there are two risk-neutral agents, the insider and the market maker. At $t=0$, the insider observes a signal that informs him of the probability distribution over the countably infinite set of possible $t=1$ states, $X=\{x_{1},x_{2},\cdots\}$. Our results apply to the finite $X$ setting as a special case. The market maker has a Bayesian prior over the possible signals. At $t=0$, there is a complete market of AD securities for $t=1$ states. [^10] The model can be adapted to the incomplete market case by partitioning the state space. In this paper, we focus on the complete market case.
@@
Suppose that the states $x_{1}<x_{2}<\cdots$ are the possible $t=1$ prices of an underlying asset. In this case, any portfolio $W(\,\cdot\,)$ of AD securities can be replicated using a combination of the underlying asset and options: [^12] See Back (2010) Exercise 3.5. Let $x_{i_{0}}$ be the expected $t=1$ asset price under the market maker’s prior at $t=0$. For any $W(\,\cdot\,)$, there exists $(a_{i})$ such that
$\!\!\!W(x_{j})$ $=a_{i_{0}}(x_{j}-x_{i_{0}})+\sum_{i<i_{0}}a_{i}(x_{j}-x_{i})_{-}+\sum_{i>i_{0}}a_{i}(x_{j}-x_{i})_{+},\quad\mbox{for all $j=1,2,\cdots$}.$ (7)
@@
The insider learns the true payoff distribution $\eta(\,\cdot\,|s_{k})$ of the underlying asset from his private signal $s_{k}$ and then submits orders for the underlying asset and options. Noise trades follow $\varepsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$ in each market $i$ (for the asset or option at strike $x_{i}$). When the insider submits orders $(a_{i})$, the order flow received by the market maker is given by
$a_{i}+\varepsilon_{i},\quad i=1,2,\cdots.$
@@ 5.1 First-Order Condition
The insider’s marginal utility from adding a marginal portfolio $v$ to $W$, $\partial\!\kern 0.80002ptJ(v;W)$, decomposes into
@@
$\underbrace{\partial\!\kern 0.80002ptJ_{p}(v;W)}_{\text{payoff of$v$}}$ $=\sum_{i}v(x_{i})\eta(x_{i}|s_{k}),$ (15)
@@ 6.1 Canonical Game
The canonical transformation (22) normalizes the trading game by the insider’s information intensity. To see the intuition, suppose the $\eta(\,\cdot\,|s_{k})$’s have non-overlapping payoffs across the state space, such as in Section 2. In this case, the information intensity matrix ${\bf L}$ is diagonal, and this transformation rescales each $\eta(\cdot|s_{k})$ by $\frac{1}{c_{k}}$, where $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$ is the total noise-adjusted payoff variation across securities conditional on $s_{k}$. In turn, this transforms ${\bf L}$ to the identity matrix—thereby normalizing the information intensity across securities and signals. [^17] Here, we note again that our proofs go through for any positive function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$ of states and signals, including the rescaled payoffs. This normalization yields an isomorphic game that is invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$ and symmetric across signals.
@@ 7.1 Informed Demand
(a)/(b): the high volatility signal/straddle, (c)/(d): low volatility signal/butterfly.
@@ 7.2 Price Impact
In Corollary 5.2, substituting $W^{*}$ for $W$ and $\widetilde{W}$ in (21) yields the equilibrium expected price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$
@@
where ${\operatorname{\mathbb{Cov}}}(\,\cdots\,,\,\cdots\,\big{|}\omega)$ is the covariance over signals $s_{k}\in S$ under the market maker’s equilibrium posterior, conditional on order flow $\omega$. The expectation $\mathbb{E}[\,\cdot\,]$ is taken over $\omega$ with respect to its equilibrium distribution. Removing this expectation in (34) yields the equilibrium price impact conditional on $\omega$.
@@
The equilibrium cross price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$ is the covariance of their payoffs, adjusted by factors related to informed and noise trades: (i), (ii), and (iii) below.
@@
Thus, $\Lambda_{i,j}$ extends the intuition of the cross price impact discussed in Section 2, showing the robustness of this basic intuition for cross-market information dynamics.
@@
    (Negative Cross Price Impact.) If security $x_{i}$ pays off only under a signal $s_{k_{i}}$ and $x_{j}$ only under $s_{k_{j}}\neq s_{k_{i}}$, their equilibrium cross price impact is negative. This is the same scenario as Example 5.3(ii), but in equilibrium. Upon receiving a buy order for $x_{i}$, the market maker infers that $x_{j}$ is unlikely to pay off and lowers its price accordingly, and vice versa. Indeed,
@@
    (Zero Cross Price Impact.) If a security $x_{j}$’s payoff has zero variation across signals (i.e., $\eta(x_{j}|\,\cdot\,)$ is constant), then its order flow has zero price impact across markets. In turn, no-arbitrage implies that informed demand for $x_{j}$ must be zero. Conversely, if the informed demand for $x_{j}$ is zero across signals, its order flow has zero information content and, therefore, zero price impact on all markets. This is exemplified by the security $x_{2}$ indicated in Figure 4c. Figures 4b and 4d show the corresponding informed demand $W^{*}(x_{2}|\,\cdot\,)=0$ to be indeed zero across signals.
@@
    (Positive Cross Price Impact.) If two securities have identical payoffs, then their cross price impact is naturally positive. This is the case for the securities $x_{1}$ and $x_{3}$ indicated in Figure 4c.
The general formula for cross price impact between derivatives is as follows.
@@
In other words, a high cross price impact between the put-call pair in a long-volatility straddle means that volatility is underpriced and should therefore correct upwards as volatility price discovery takes place. These insights extend to higher moments beyond volatility, offering testable empirical predictions (see Section 8).

### v28

@@ Abstract
Keywords: Price Discovery, Contingent Claims, Options, Straddle, Volatility Smile.
@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information about an asset’s expected payoff, that information is incorporated into the asset price. More broadly, agents often have information about an asset’s payoff beyond its mean, such as volatility, and to leverage such information they must trade derivatives. Indeed, a principal utility of derivatives is to offer investors targeted exposure to various aspects of the underlying asset’s payoff distribution. Such derivatives trading has long been standard practice, as covered in popular textbooks such as Hull (2015). This implies that a wide range of information beyond just the expected payoff is incorporated into derivatives prices.
Yet, despite their practical importance, the existing literature lacks a basic model that articulates how derivative markets systematically incorporate diverse information signals into prices. This gap stems largely from the challenges of modeling the complex information dynamics across derivatives. We address this gap by developing a model that fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. Integrating the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985), we obtain general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
@@
Replicating derivatives, such as options, using AD securities provides equivalent formulations of our model that account for the structural linkages across derivatives. In an equivalent options formulation, the informed trader has information about an underlying asset’s payoff distribution and trades options, with no restriction on the payoff distribution or his information.
@@
The informed demand formula we obtain is simple and practical. It constructs the informed trader’s contingent-claim portfolio by assigning positive portfolio weight to the true payoff distribution and negative weights to the others. Translated to options, the formula prescribes option strategies for trading on any given aspect of the underlying payoff distribution. In particular, it offers the first equilibrium rationalization for widely used option trading strategies.
@@
When trading is restricted to only the underlying asset, our model reduces to the single-asset price discovery framework of Kyle (1985), where the equilibrium supply function is linear. The slope of the inverse supply function, known as “Kyle’s lambda,” determines the price impact and the extent to which demand reveals information about the asset’s payoff. This slope is proportional to the variation of the asset’s payoff. In our more general AD setting, demand for one security impacts prices across all markets based on the information it conveys about other securities. We show that the equilibrium price impact between two securities is proportional to their payoff covariance, generalizing Kyle’s lambda to an infinite-dimensional matrix of payoff covariances.
@@
Across option markets, trades in one option often impact the prices of others as options trades jointly convey information about the underlying payoff. For example, a common volatility trade involves trading a put-call pair to profit from price movements regardless of direction. Therefore, trades in such pairs jointly convey volatility information. Our characterization of such cross-market information spillovers generates new empirical predictions linking cross-option dynamics to various aspects of the underlying payoff. Moreover, we offer an equilibrium explanation for the “volatility smile,” an empirical regularity of option-implied volatilities that has been widely studied but only in reduced-form.
@@
The joint supply schedules for all AD securities form a pricing kernel, under asymmetric information. In contrast to the one-dimensional linear supply function for a single asset, this pricing kernel is a high-dimensional nonlinear object that prices all derivatives simultaneously. While the pricing of derivatives has been extensively studied from the hedging perspective since Black and Scholes (1973), we go beyond hedging under complete information, providing a microfoundation for derivatives trading and price patterns that hedging motives alone cannot explain.
@@
While empirical evidence (e.g., Roll et al. (2010)) suggests derivatives improve market efficiency, theoretical explanations are limited. Our model offers a clear explanation: the linear equilibrium of Kyle (1985) is not robust with respect to the introduction of derivatives, as the absence of these markets constrains the informed trader to suboptimal linear demands. Derivatives allow the informed trader to optimally reduce his demand for low-probability states (see Example 7.1). By incorporating information that cannot be reflected in linear trades of the underlying asset, derivatives improve market efficiency.
@@
In a complete market with derivatives, our model demonstrates that the information efficiency of all prices is summarized by a single measure. Notably, this measure remains invariant to variations in the possible payoff distributions or noise trading levels across markets. Thus, the informativeness of derivatives prices is robust to specific information details.
@@
Kyle (1985) provides the seminal framework for single-asset price discovery. Extensions by Caballe and Krishnan (1994), Foster and Viswanathan (1996), and Back et al. (2000) consider settings with multiple insiders holding differential information across assets. Rochet and Vila (1994) relax the assumption of normal payoff distributions. We expand the scope of this literature to encompass joint price discovery across contingent claim markets.
@@
Empirical research has consistently shown that trades and prices across options markets reflect information beyond the expected payoff. Pan and Poteshman (2006) and Ni et al. (2008) find that option trading volumes predict future stock returns and volatility, while Goyal and Saretto (2009) show that volatility mispricing can be exploited through option straddles. Other studies highlight how option order flows can signal firm value and market expectations of future volatility—see Roll et al. (2009), Augustin et al. (2019), and Cao et al. (2005). Despite substantial empirical evidence of joint price formation across option markets, existing theoretical models remain limited to no-option versus single-option comparisons (e.g., Back (1993); Biais and Hillion (1994); Easley et al. (1998); Collin-Dufresne et al. (2021)). This leaves an empirical-theoretical gap in understanding how trades and prices across options jointly reflect broader information. Our model bridges this gap by directly capturing the richness of information embedded across option trades and prices. It provides a unified theoretical framework that consolidates empirical findings and suggests new avenues of empirical research.
@@
Different from our strategic setting, there have also been theoretical studies on non-strategic informed trading across multiple securities, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). These non-strategic models do not explain the common derivatives trading strategies or observed cross-strike empirical regularities. Moreover, Admati (1985) and Chabakauri et al. (2022) impose restrictive parametric assumptions on the payoff distribution, and Malamud (2015) depends on properties intrinsic to the continuum. We make no parametric assumptions on the payoff, and our results do not depend on the discrete or continuum modeling choice. Additionally, we derive testable predictions linking the underlying payoff characteristics, such as higher moments, to observed price impacts across options. This new insight is unique to the strategic setting.
@@
Many option pricing and financial econometric models have sought to capture the volatility smile by correcting misspecifications in the Black and Scholes (1973) framework. Prominent classes of such models include the jump-diffusion, [^1] See Merton (1976) and Bates (1991). stochastic volatility, [^2] See Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), and Aït-Sahalia et al. (2021). and local volatility models. [^3] See Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012). Our framework moves beyond reduced-form model misspecifications to directly explain the strike structure of implied volatility via cross-strike information dynamics.
@@
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results through a simplified setting. Section 3 presents the general model. Section 4 identifies the sufficient statistic for the uninformed trader’s cross-market inferences. Section 5 analyzes the informed trader’s cross-market price impact. Section 6 establishes equilibrium through a canonical reduction of the general trading game. While this canonical reduction is a key modeling innovation that makes our general model tractable, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be understood from Example 6.7 without delving into the precise details. Section 7 discusses the characterization of price discovery across contingent claims—the informed demand, price impact, and information efficiency of prices. Section 8 explores the immediate empirical implications of our results for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@ 2 Basic Intuition
We begin with a simplified setting that encapsulates the basic economic intuition we bring out. In the general model, we will develop the robustness of this intuition—without restriction on the number of securities, payoff distributions, or type of private information.
@@ 3 General Model
As in Section 2, there are two risk-neutral agents, the insider and the market maker. At $t=0$, the insider observes a signal that informs him of the probability distribution over the set of possible $t=1$ states, $X=\{x_{1},x_{2},\cdots\}$ (which is countably infinite, with finite $X$ included as a special case). The market maker has a Bayesian prior over the possible signals. At $t=0$, there is a complete market of AD securities for $t=1$ states. [^10] The model can be adapted to the incomplete market case by partitioning the state space. In this paper, we focus on the complete market case. There is a risk-free asset in perfectly elastic supply at risk-free rate zero.
@@
Suppose the states $x_{1}<x_{2}<\cdots$ are the possible $t=1$ prices of an underlying asset. In this case, any portfolio $W(\,\cdot\,)$ of AD securities can be replicated using a combination of the underlying asset and options: [^12] See Back (2010) Exercise 3.5. Let $x_{i_{0}}$ be the expected $t=1$ asset price under the market maker’s prior at $t=0$. For any $W(\,\cdot\,)$, we can find $(a_{i})$ such that
$\!\!\!W(x_{j})$ $=a_{i_{0}}(x_{j}-x_{i_{0}})+\sum_{i<i_{0}}a_{i}(x_{j}-x_{i})_{-}+\sum_{i>i_{0}}a_{i}(x_{j}-x_{i})_{+},\quad\mbox{for $j=1,2,\cdots$}.$ (7)
@@
The insider learns the true payoff distribution $\eta(\,\cdot\,|s_{k})$ of the underlying asset from his private signal $s_{k}$ and then submits orders for the underlying asset and options. Noise trades follow $\varepsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$ in each market $i$ (for the asset or option at strike $x_{i}$). When the insider submits orders $(a_{i})$, the order flow received by the market maker is given by $a_{i}+\varepsilon_{i}$, for $i=1,2,\cdots.$
@@ 5.1 First-Order Condition
The insider’s marginal utility from adding a marginal portfolio $v$ to a portfolio $W$ decomposes into
@@
$\underbrace{\partial\!\kern 0.80002ptJ_{p}(v;W)}_{\text{marginal payoff}}$ $=\sum_{i}v(x_{i})\eta(x_{i}|s_{k}),$ (15)
@@ 6.1 Canonical Game
The canonical transformation (22) normalizes the trading game by the insider’s information intensity. For intuition, suppose the $\eta(\,\cdot\,|s_{k})$’s have non-overlapping payoffs across states, such as in Section 2. In this case, there is zero covariation of payoffs across signals, which results in a diagonal information intensity matrix ${\bf L}$. The $k$-th diagonal entry of ${\bf L}$ is $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$, which is the noise-adjusted payoff variation across securities conditional on $s_{k}$. The canonical transformation rescales each $\eta(\cdot|s_{k})$ by $\frac{1}{c_{k}}$, i.e., it normalizes the payoff variations to one across signals, thereby transforming ${\bf L}$ to the identity matrix. [^17] We note again that our proofs apply to any positive function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the rescaled payoffs. This normalization yields an isomorphic game that is invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$ and symmetric across signals.
@@ 7.1 Informed Demand
(a)/(b): high volatility signal/straddle, (c)/(d): low volatility signal/butterfly.
@@ 7.2 Price Impact
In Corollary 5.2, substituting $W^{*}$ for $W$ and $\widetilde{W}$ in (21) yields the equilibrium expected price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$:
@@
where ${\operatorname{\mathbb{Cov}}}(\,\cdots\,,\,\cdots\,\big{|}\omega)$ is the covariance over signals $s_{k}\in S$ under the market maker’s equilibrium posterior, conditional on order flow $\omega$. The expectation $\mathbb{E}[\,\cdot\,]$ is taken over $\omega$ with respect to its equilibrium distribution. By removing this expectation in (34), we obtain the equilibrium price impact conditional on order flow $\omega$.
@@
The equilibrium cross price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$ is determined by their payoff covariance, adjusted for factors related to the informed and noise trades:
@@
Thus, $\Lambda_{i,j}$ extends the intuition behind cross-market price impact discussed in Section 2.
@@
    (Negative Cross Price Impact.) If security $x_{i}$ pays off only under signal $s_{k_{i}}$ and $x_{j}$ only under $s_{k_{j}}\neq s_{k_{i}}$, their equilibrium cross price impact is negative. This is the same scenario as Example 5.3(ii), but in equilibrium. Upon receiving a buy order for $x_{i}$, the market maker infers that $x_{j}$ is unlikely to pay off and lowers its price accordingly, and vice versa. Indeed,
@@
    (Zero Cross Price Impact.) If a security $x_{j}$’s payoff has zero variation across signals (i.e., $\eta(x_{j}|\,\cdot\,)$ is constant), then its order flow has zero price impact across markets. In turn, no-arbitrage implies that informed demand for $x_{j}$ must be zero. Conversely, if the informed demand for $x_{j}$ is zero across signals, its order flow is pure noise and, therefore, has zero price impact across markets. This is illustrated by the security $x_{2}$ shown in Figure 4c. Figures 4b and 4d confirm the corresponding informed demand $W^{*}(x_{2}|\,\cdot\,)=0$ to be indeed zero across signals.
@@
    (Positive Cross Price Impact.) If two securities have identical payoffs, then their cross price impact is naturally positive. This is illustrated by the securities $x_{1}$ and $x_{3}$ shown in Figure 4c.
We obtain the following general formula for cross price impact between derivatives.
@@
In other words, a high cross price impact between the put-call pair in a long-volatility straddle means that volatility is underpriced and should therefore correct upwards as volatility price discovery takes place. These insights regarding cross-market information dynamics generalize to higher moments beyond volatility, offering testable empirical predictions in a broad framework (see Section 8).